OpenStudy (anonymous):

determine whether the ordered pair is a solution of the inequality (2,10); 6y-7x>5

OpenStudy (anonymous):

6(10) - 7(2) > 5 60 - 14 > 5 26 > 5 yes it is a solution.

jimthompson5910 (jim_thompson5910):

6y-7x>5 6(10)-7(2)>5 60-14>5 46>5 Since the last inequality is true, this means that (2,10) is a solution

OpenStudy (anonymous):

sorry lol hit 2 instead of 4, getting a little sleepy :D

OpenStudy (anonymous):

\[\left(\begin{matrix}2x-5 \\ 8\end{matrix}\right)\le-4\]

jimthompson5910 (jim_thompson5910):

\[\large \frac{2x-5}{8}\le -4\] \[\large 2x-5\le -4*8\] \[\large 2x-5\le -32\] \[\large 2x\le -32+5\] \[\large 2x\le -27\] \[\large x\le -\frac{27}{2}\]